OpenStudy (anonymous):

Help! Fan/medal The Tran family and the James family each used their sprinklers last summer. The water output rate for the Tran family's sprinkler was 30L per hour. The water output rate for the James family's sprinkler was 15L per hour. The families used their sprinklers for a combined total of 50 hours, resulting in a total water output of 1200 L. How long was each sprinkler used?

OpenStudy (anonymous):

30T + 15J = 1200 T + J = 50

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

isolate T in the second equation

OpenStudy (anonymous):

do i add from here?

OpenStudy (anonymous):

the equations?

OpenStudy (anonymous):

nope. T + J = 50 how do you get T by itself?

OpenStudy (anonymous):

I ended up getting T=30 and J=60